OpenStudy (anonymous):

[6.02] and [6.03] Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 16 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

x-y=16 2x+y=2 y=-2x+2 x-(-2x+2)=16 x+2x-2=16 3x=18 x=6 y=-10 i used substitution because it was easy to isolate the y variable.

OpenStudy (anonymous):

1. x - y = 16 ----> x = y + 16 2x + y = 2 Now substitute y + 16 in for x in the 2nd equation... 2(y + 16) + y = 2 ( multiply through the parenthesis) 2y + 32 + y = 2 (combine like terms) 3y + 32 = 2 (subtract 32 from both sides to get 3y by itself) 3y + 32 - 32 = 2 - 32 3y = 2 - 32 3y = - 30 (divide by 3 to get y by itself) 3y/3 = - 30/3 y = - 10 Now substitute - 10 in for y in 1st equation to find x x - y = 16 x - (- 10) = 16 x + 10 = 16 (subtract 10 from each side to get x by itself) x + 10 - 10 = 16 - 10 x = 16 - 10 x = 6 Now we can check our answers...y = - 10 and x = 6 x - y = 16 10 - (-6) = 16 10 + 6 = 16 16 = 16 (correct) You can also check it with the other equation... 2x + y = 2 2(6) + (- 10) = 2 12 - 10 = 2 2 = 2 (correct) ANSWER : x = 6 and y = - 10 2. Substitution is easier to use then addition/elimination when there is only two equations.